# On the Vulnerability of Concept Erasure in Diffusion Models

### v2

@@ Appendix A Choice of loss function
where $z_{t}$ is obtained through the forward diffusion process, ${z_{0}\sim p_{\text{target}}}$ is sampled from the target data distribution, $y_{\text{target}}$ is the target prompt. $L_{1}$ [27] and $L_{2}$ [3] are minimized by optimizing prompts $y$ to match the denoiser predictions respectively to: the true noise from the forward diffusion sequence $\epsilon$, or the predicted noise by the baseline denoiser with the target prompt $\epsilon_{\theta}\big{(}z_{t},t,\mathcal{T}(y_{\text{target}})$.
To compare both loss functions, we consider the non-seed-specific prompt embedding attack setting:
@@
When carrying about embedding attacks on AdvUnlearn, which modifies only the text encoder rather than the denoiser, embeddings which can successfully restore concepts are guaranteed to exist. It is thus surprising to observe that in this seemingly easier setting, the ASR on AdvUnlearn is not necessarily higher than other denoiser-based concept erasure methods.
The RECORD algorithm discussed in this paper uses loss (4) by default due to its marginal improvement in the attack performance. Consequently, this particular loss may result in the increased runtime of the RECORD algorithm, as well as requiring access to the baseline model $\epsilon_{\theta}$. These limitations, however, can be mitigated or avoided by switching to loss (3) at the expense of a marginally poorer performance.
@@
Table 5: Attack success rate (%) of the embedding attacks for art style, object and nudity attacks. Embed attack 1 and 2 refer to the two loss formulation $\widetilde{L}_{1}$ and $\widetilde{L}_{2}$, respectively.
Appendix B Example Images
In this section we present example images generated by different restoration methods, which extends from Table 1, as well as from the embedding attack discussion in Appendix A
Erased Concept Restoration Method Erasure Method
@@ Appendix B Example Images
[10] [45] [20] [25] [11] [46]
Table 6: Generated images of erased concepts using embedding attacks. Each image column of the same concept is generated using the same latent initialization. Models have unlearned the style of van gogh paintings. Target prompt "a painting in the style of van gogh"
Erased Concept Restoration Method Erasure Method
ESD SH SPM SalUn UCE AdvUnlearn
[10] [42] [25] [8] [11] [46]
Table 7: Generated images of erased concepts using embedding attacks. Each image column of the same concept is generated using the same latent initialization. Models have unlearned the concept of nudity. Target prompt asked for nudity.
Erased Concept Restoration Method Erasure Method
ESD ED SH SPM SalUn AdvUnlearn
[10] [43] [42] [25] [8] [46]
Table 8: The generated images of erased concepts using embedding attacks. Each image column of the same concept is generated using the same latent initialization. Target prompt "a photorealistic image of object"
Erased Concept Restoration Method Erasure Method
ESD ED SH SPM SalUn AdvUnlearn
[10] [43] [42] [25] [8] [46]
Church P4D [3] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD [47] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Garbage Truck P4D [3] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD [47] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Parachute P4D [3] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD [47] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 9: The generated images of erased concepts using prompt attacks. Each image column of the same concept is generated using the same latent initialization. Target prompt "a photorealistic image of object"
Table 10: The generated images of erased concepts using prompt attacks. Each image column of the same concept is generated using the same latent initialization. Target prompt asked for nudity.
Erased Concept Restoration Method Erasure Method
ESD ED SH SPM SalUn AdvUnlearn
[10] [43] [42] [25] [8] [46]
Nudity P4D [3] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD [47] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix C Tokens Length Ablation Study
The number of available tokens for perturbation is an important factor affecting the ASRs of the concept restoration methods. This section showcases the ASR of RECORD with 16 and 32 tokens, as shown in Tables 11 and 12. The rest of the optimization configuration is consistent with that of Section 4.2.
Erased Concept Prompt Length Erasure Method
ESD FMN AC SPM UCE AdvUnlearn
[10] [45] [20] [25] [11] [46]
@@ Appendix C Tokens Length Ablation Study
Table 11: Attack success rate (%) for models with the van Gogh style erased. Comparing RECORD with different allowed prompt lengths in terms of tokens.
Erased Concept Prompt Length Erasure Method
@@
Table 12: Attack success rate (%) for models with the concepts Church, Garbage Truck, Parachute, and Nudity erased. Comparing RECORD with different allowed prompt lengths in terms of tokens.

### v3

@@ Appendix B Embedding-level Attack Results and Loss Function Comparison
where $z_{t}$ is obtained through the forward diffusion process, $z_{0}\sim p_{\text{target}}$} is sampled from the target data distribution, $y_{\text{target}}$ is the target prompt. $L_{1}$ Pham et al. (2023) and $L_{2}$ Chin et al. (2023) are minimized by optimizing prompts $y$ to match the denoiser predictions respectively to: the true noise from the forward diffusion sequence $\epsilon$, or the predicted noise by the baseline denoiser with the target prompt $\epsilon_{\theta}\big(z_{t},t,\mathcal{T}(y_{\text{target}}))$.
To compare both loss functions, we consider the embedding-level attack setting:
@@
Table 4: Attack success rate (%) of the embedding-level attacks for art style, object and nudity attacks. Embed attack 1 and 2 refer to the two loss formulation $\widetilde{L}_{1}$ and $\widetilde{L}_{2}$, respectively.
The RECORD algorithm discussed in this paper uses loss Equation [4] by default due to its marginal improvement in the attack performance. Consequently, this particular loss may result in the increased runtime of the RECORD algorithm, as well as requiring access to the baseline model $\epsilon_{\theta}$. These limitations, however, can be mitigated or avoided by switching to loss Equation [3] at the expense of a marginally poorer performance.
Appendix C Ablation Study on Tokens Lengths
This section presents an ablation study to examine the effect of two key parameters of RECORD: the number of tokens available for perturbation $S$ and number of passes on the token sequence $N$. The product of these two parameters, $N\times S$, corresponds to the total number of optimization steps, which linearly scales the algorithm’s runtime. The rest of the parameter configuration is consistent with that of Section 4.2.
Tables 5 and 6 demonstrate a strong impact of both $N$ and $S$ on RECORD’s performance. As expected, increasing these parameters generally leads to better performance. However, a notable finding is that even with $S=16$, RECORD still achieves decent ASRs that often surpass those of P4D and UD with $S=64$ in Table 3. For $S=16$, RECORD also has a much shorter runtime as shown in Figure 3.
We also observed that reducing $S$ has a more pronounced negative effect on ASR than reducing $N$. This is evident in Table 6 (right), where a configuration with a smaller $S$ but a compensating increase in $N$, for maintaining a consistent total number of optimization steps, still resulted in a decay in ASR. This decay, however, was less severe than a simple reduction in $S$ in Table 5 without a compensatory increase in $N$. This suggests the breadth of the search space of the adversarial tokens is more critical than the depth of the search, and choosing a larger $S$ is the preferred strategy for maximizing performance for the same compute cost.
Erased Concept Prompt Length $S$ Erasure Method
@@ Appendix C Ablation Study on Tokens Lengths
2023a 2024a 2023 2024 2023b 2024b
@@
Erased Concept Prompt Length $S$ Erasure Method
@@
Table 5: Attack success rate (%) of RECORD with different prompt length $S$.
Number of Passes NN Erasure Method
ESD AC AdvUnlearn
2023a 2023 2024b
5 50.8 91.0 7.8
10 58.4 92.4 28.2
15 65.8 93.6 22.0
20 64.0 94.0 33.0
Number of Passes NN Prompt Length $S$ Erasure Method
ESD AC AdvUnlearn
2023a 2023 2024b
80 16 33.6 91.4 30.4
40 32 45.8 91.2 33.8
20 64 64.0 94.0 33.0
Table 6: Attack success rate (%) of RECORD with different number of passes NN (left) and with fixed number of optimization steps (right) on models unlearned on van Gogh style.
Appendix D Ablation Study on Gradient and Candidate Tokens
In this section, we conduct ablation studies on the gradient token number $J$ and candidate token number $K$ to justify the design choices as well as the performance and efficiency of the RECORD algorithm. The rest of the experimental setting is consistent to that of Section 4.2.
Our experiments demonstrate that increasing the number of gradient tokens $J$, which are used for linearised gradient estimation, yields only marginal performance improvements in terms of ASRs. In contrast, the number of candidate tokens $K$ for exact evaluation has a more significant impact on boosting ASRs.
This finding suggests it is possible to substantially accelerate the algorithm by significantly reducing $J$ to accelerate the algorithm. Notably, with $J=8$, the runtime of the algorithm can be reduced by 60%, while having only a marginal relative drop in ASRs of $5-10\%$, or $3-4\%$ in absolute terms. Smaller $J$ also helps in reducing memory consumption, as gradient estimation through backpropagation with $J$ tokens corresponds to a major but now mitigable memory bottleneck.
Gradient Token $J$ Erasure Method
ESD AC AdvUnlearn
2023a 2023 2024b
1 60.2 93.8 16.4
8 61.4 94.6 30.2
16 64.0 92.2 30.0
32 65.2 92.8 30.6
64 64.0 94.0 33.0
Candidate Token $K$ Erasure Method
ESD AC AdvUnlearn
2023a 2023 2024b
1 17.4 65.8 2.4
8 54.6 93.0 8.4
16 54.2 90.4 17.8
32 61.2 95.2 29.8
64 64.0 94.0 33.0
Table 7: Combined tables showing van Gogh Gradient and Candidate Tokens.
Prompt Length $S$ Restoration Method Runtime/s/s
P4D UD RECORD (J=64) RECORD (J=8)
8 2059±49 305±7 235±16 97±2
16 2050±34 429±7 464±18 191±2
32 2070±22 638±39 929±16 382±5
64 2079±22 1003±50 1859±15 832±7
Table 8: The mean and the standard deviation of the restoration method runtime are computed over 10 runs.
Appendix E Fixed Seed Attacks
To assess the performance of RECORD in the fixed-seed setting, we follow the experimental setup of UD Zhang et al. (2023), where the target prompt-seed pairs are taken from the I2P dataset Schramowski et al. (2023), which fixes the generation seed. In these experiments, we also similarly follow the previous works of using a NudeNet classifier Bedapudi (2019) and evaluates the optimized prompts on-the-fly: if a generated image is deemed unsafe, the optimization stops immediately.
Under this setting, we note RECORD achieves highly competitive performance in comparison with P4D and UD, especially on concept-erased models that are easier to attack, such as ESD, SalUn, and AdvUnlearn.
Erased Concept Restoration Method Erasure Method
ED ESD SalUn SH SPM AdvUnlearn
2024 2023a 2023 2024 2024 2024b
Nudity P4D 2023 3.4 84.8 21.2 4.2 100.0 27.1
UD 2023 1.7 87.3 19.5 11.0 100.0 22.0
RECORD 1.7 98.3 30.5 9.3 100.0 41.5
Table 9: Attack success rate (%) of concept restoration methods in the fixed-seed setting.
Appendix F Adversarial Prompt Transferability
This section investigates the transferability of adversarial prompts generated by different concept erasure and restoration methods. Specifically, we examine whether prompts optimized on an unlearned SD1.4 model can successfully generate erased concepts on other models, such as SDXL and FLUX unlearned using ESD. This study assesses the generalizability and robustness of these prompts across different model architectures and explores the feasibility for an external adversary to use prompts optimized on a white-box open-source model to attack a different black-box model.
Our analysis uses a collection of adversarial prompts that can successfully generate erased concepts on their correspondingly unlearned SD1.4 models. Any prompts that fail to generate the erased concepts are not included in this study. As shown in Table 10, the adversarial prompts identified by the RECORD method exhibit greater transferability than those from P4D and UD, often by a significant margin. This suggests that the optimization strategy of RECORD produces prompts that are more robust and less model-specific.
We also observed that prompt transferability is highly dependent on both the specific erased concept and the target model. Adversarial prompts from SD1.4 generally transferred better to SDXL than to FLUX. This is likely due to the significant differences in training data and model architecture between SD1.4 and FLUX. Despite these differences, it is interesting to note that some level of transferability can still be preserved. This suggests a fundamental, underlying transferability of adversarial prompts in general, indicating that they retain some semantic meaning even when they appear as human-unreadable, gibberish-like strings.
Erased Concept Model ESD AC AdvUnlearn
P4D UD RECORD P4D UD RECORD P4D UD RECORD
van Gogh SDXL 2.0 4.0 5.0 4.4 3.8 6.0 1.6 3.4 10.8
FLUX 0.2 0.0 1.4 0.0 0.4 1.8 0.8 0.2 2.4
Erased Concept Model ESD SH AdvUnlearn
P4D UD RECORD P4D UD RECORD P4D UD RECORD
Church SDXL 64.4 58.6 79.8 32.0 52.6 59.8 49.2 64.6 61.6
FLUX 5.2 2.6 14.8 6.2 0.2 8.4 1.0 1.0 5.4
Nudity SDXL 61.8 68.0 77.6 11.8 75.4 22.0 49.8 74.2 66.0
FLUX 18.0 18.4 38.0 5.4 16.4 9.6 8.2 11.6 22.4
Table 10: Attack success rate (%) of using successful adversarial prompts on SD1.4 on ESD-unlearned SDXL and FLUX.
Appendix G Scalability on Large Models
We extend RECORD to work on larger models, namely SDXL Podell et al. (2024) and FLUX Labs et al. (2025). The major challenge is that SDXL and FLUX both use two separate text encoders for encoding input texts. This differs from SD1.4, where only one text encoder is used in its pipeline. This dual-encoder setup leads to difficulties in the candidate selection stage. For SDXL, both CLIP text encoders share a similar tokenizer with consistent token_id mapping, but with different embedding tables $E_{1}$, $E_{2}$ and different CLIP encoders. In this case, we propose three different strategies for handling candidate selection:
  • Random Switching: At each optimization step, randomly choose one of the two text encoders for computing score and select the top-$K$ tokens.
    $\mathcal{V}=\text{Top}_{K}(E_{i}g_{i}),\quad\text{where }i\sim\text{Uniform}\{1,2\},$
    where $g_{i}$ is the gradient of the corresponding embedding table $E_{i}$.
  • Interleaving: Compute two separate scores and select the top-$K/2$ tokens with respect to each text encoder. In our experience, the number of overlapping candidate tokens are negligible compared to the size of $K$.
    $\mathcal{V}=\text{Top}_{K/2}(E_{1}g_{1})\cup\text{Top}_{K/2}(E_{2}g_{2}).$
  • Blend: Compute a mixture of the scores from both text encoders and select the top-$K$ tokens.
    $\mathcal{V}=\text{Top}_{K}\Big((1-\alpha)E_{1}g_{1}+\alpha E_{2}g_{2}\Big),$
    where $\alpha\in[0,1]$ is a tunable hyperparameter. When $\alpha=0$ or $1$, this is equivalent to optimizing over only one text encoder.
We denote the embedding tables of CLIP-ViT/L (CLIPTextModel) and CLIP-ViT/G (CLIPTextModelWithProjection) as $E_{1}$ and $E_{2}$, respectively. The ASRs of the three strategies are presented in Table 11. We note that, for SDXL, using interleaving strategy or optimizing only on CLIP-ViT/L works the best. Example images can be found in Appendix K.
FLUX uses CLIP and T5 Raffel et al. (2019) as its text encoders with two completely different tokenizers. Tokens of the two text encoders thus correspond to different strings of text. This is exacerbated by the non-bijective nature of the token-string mappings, i.e. original token $\to$ string $\to$ token $\neq$ original token. These make Interleaving and Blend not applicable, unless $\alpha\in\{0,1\}$. Here we consider the embedding tables of CLIP and T5 encoder as $E_{1}$ and $E_{2}$. RECORD performs best on FLUX when only optimizing on the CLIP encoder (Table 12), which is mostly consistent with the behavior in SDXL.
Candidate Selection Strategy Erased Concepts Average
van Gogh Church Nudity
Random switching 17.4 86.0 23.6 32.5
Interleaving 26.6 92.4 17.6 34.8
Blend $\alpha=0.0$ 19.0 92.4 26.6 35.4
Blend $\alpha=0.5$ 14.2 91.0 17.6 31.4
Blend $\alpha=1.0$ 7.2 89.2 15.0 28.4
Table 11: Attack success rate (%) of different candidate selection strategy for SDXL.
Candidate Selection Strategy Erased Concepts Average
van Gogh Church Nudity
Random switching 1.6 19.8 8.2 9.8
Blend $\alpha=0.0$ 5.0 54.6 14.4 24.7
Blend $\alpha=1.0$ 2.6 0.2 1.4 1.4
Table 12: Attack success rate (%) of different candidate selection strategy for FLUX.
Appendix H LLM Usage Declaration
Large language models have been occasionally used in this project for polishing writing, suggesting and applying bug fixes with significant human oversight, and for interfacing with computing infrastructure such as Slurm and Kubernetes.
Appendix I Embedding-level Attack Example Images
In this section we present SD1.4 example images generated by different restoration methods, as well as from the embedding-level attack discussion in Appendix B
Erased Concept Restoration Method Erasure Method
ESD FMN AC SPM UCE AdvUnlearn
Gandikota et al., (2023a) Zhang et al., (2024a) Kumari et al., (2023) Lyu et al., (2024) Gandikota et al., (2023b) Zhang et al., (2024b)
Table 13: Generated images of erased concepts using embedding-level attacks. Each image column of the same concept is generated using the same latent initialization. Models have unlearned the style of van Gogh paintings. Target prompt "a painting in the style of van Gogh"
Erased Concept Restoration Method Erasure Method
ESD SH SPM SalUn UCE AdvUnlearn
2023a 2024 2024 2023 2023b 2024b
Table 14: Generated images of erased concepts using embedding-level attacks. Each image column of the same concept is generated using the same latent initialization. Models have unlearned the concept of nudity. Target prompt asked for nudity.
Appendix J Token-level Attack Example Images
Erased Concept Restoration Method Erasure Method
ESD ED SH SPM SalUn AdvUnlearn
2023a 2024 2024 2024 2023 2024b
Table 15: The generated images of erased concepts using embedding-level attacks. Each image column of the same concept is generated using the same latent initialization. Target prompt "a photorealistic image of {object}"
Erased Concept Restoration Method Erasure Method
ESD ED SH SPM SalUn AdvUnlearn
2023a 2024 2024 2024 2023 2024b
Church P4D 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Garbage Truck P4D 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Parachute P4D 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 16: The generated images of erased concepts using token-level attacks. Each image column of the same concept is generated using the same latent initialization. Target prompt "a photorealistic image of {object}"
Table 17: The generated images of erased concepts using token-level attacks. Each image column of the same concept is generated using the same latent initialization. Target prompt asked for nudity.
Erased Concept Restoration Method Erasure Method
ESD ED SH SPM SalUn AdvUnlearn
2023a 2024 2024 2024 2023 2024b
Nudity P4D 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix K Example Images from Larger Models
Van Gogh Church Nudity
No attack [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
RECORD [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 18: Images generated by ESD erased FLUX on van Gogh paintings, churches and nudity.
Van Gogh Church Nudity
No attack [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
RECORD [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 19: Images generated by ESD erased SDXL on van Gogh paintings, churches and nudity.